OpenStudy (anonymous):

g(x) = x^4lnx what are the critical points of g?

OpenStudy (anonymous):

Use Product rule to find derivative

OpenStudy (anonymous):

Find out where your derivative function hit zero

OpenStudy (anonymous):

take log on both side. log(y)=4ln(x)log(x) (1/y)dy/dx=(4/x)log(x)+(4/x)log(x) dy/dx=y[(8/x)log(x)] x^4lnx[(8/x)log(x)]=0 from here we find either x=0 or x=1 but x cant be 0 because domain of log(x) is greater then 0. so ans is 1